# Skywork R1V2: Multimodal Hybrid Reinforcement Learning for Reasoning

### v1

@@ 4.1.1 Text Reasoning Benchmarks
  • AIME 2024: This benchmark includes competition problems from the 2024 American Invitational Mathematics Examination (AIME), a prestigious and highly selective contest for elite high school students. It assesses advanced mathematical competencies, requiring deep conceptual understanding and rigorous logical reasoning skills.
@@
  • IFEVAL [32]: A benchmark designed to evaluate conditional reasoning in large language models, focusing on the models’ ability to correctly interpret and follow "if-then" statements. It tests for both factual accuracy and logical consistency in responding to hypothetical scenarios.
@@ 4.1.2 Multimodal Reasoning Benchmarks
Baselines We conduct comprehensive evaluations against several strong proprietary models, including Claude-3.5-Sonnet (20241022) [1], OpenAI-o4-mini [18], OpenAI-o1 [9], Gemini 2 Flash [22], and Kimi k1.5 longcot [24]. Additionally, we compare our method with advanced open-source models, such as Skywork-R1V1 [19], InternVL3-38B [33], QvQ-Preview-72B [25], Deepseek R1 [7] and Qwen2.5-VL-72B-Instruct [4].
@@ 4.2 Main Results
The improvement over R1V1 is substantial, with a 5.6 percentage point increase on MMMU and a 7.0 point increase on MathVista. More impressively, R1V2 outperforms even larger models such as Qwen2.5-VL-72B (70.2% vs. 73.6% on MMMU) and QvQ-Preview-72B (70.3% vs. 73.6% on MMMU).
@@
Model MMMU Math- Olympiad AIME LiveCode live IFEVAL
Vista Bench 24 bench Bench
@@
Skywork-R1V2 (Ours) 73.6 74.0 62.6 78.9 63.6 73.2 82.9
Skywork-R1V1 68.0 67.0 - 72.0 57.2 54.6 72.5
DeepseekR1-671B - - - 79.8 65.9 71.6 83.3
InternVL3-38B 70.1 72.3 - - - - -
Qwen2.5-VL-72B 70.2 74.8 40.4 - - - -
QvQ-Preview-72B 70.3 71.4 33.2 - - - -
Proprietary Models
Claude-3.5-Sonnet 70.4 67.7 - - - - -
Gemini-2-Flash 70.7 73.1 - - - - -
Kimi-k1.5-longcot 70.0 74.9 - - - - -
OpenAI-o1 - - - 74.3 63.4 72.18 -
OpenAI-o4-mini 81.6 84.3 - 93.4 74.6 78.13 -
@@
While R1V2 is optimized for reasoning capabilities, we also evaluated its performance on general vision understanding tasks to assess potential trade-offs. In document understanding, R1V2 maintains competitive performance (81.3% on AI2D, 79.0% on ChartQA and TextVQA) despite its reasoning focus, showing only a modest gap compared to specialized models. For video understanding, R1V2 achieves acceptable results (60.2% on Video-MME, 61.5% on MVBench) but demonstrates stronger capabilities when temporal reasoning is required (1.92 on MMBench-Video compared to InternVL2.5-38B’s 1.82).
Our analysis of hallucination tendencies revealed an important finding: while aggressive optimization for reasoning initially increased hallucination rates, our MPO approach effectively mitigated this issue, reducing hallucination from 18.4% with standard SFT to 8.7% with our full implementation. This balance allows R1V2 to maintain reasonable factual accuracy (68.9% on RealWorldQA) while excelling at complex reasoning tasks. These results highlight the inherent trade-off between specialized reasoning and general visual understanding capabilities, with our current design deliberately prioritizing reasoning performance. Future versions will focus on enhancing general visual capabilities while preserving the strong reasoning foundation established in R1V2.
@@ 4.3 Qualitative Analysis
Figure 5 demonstrates R1V2’s physics reasoning capabilities through its analysis of an electromagnetic problem involving rotating magnetic fields and AC generators. When presented with a diagram showing two coils with different numbers of turns positioned perpendicular to each other, the model methodically evaluates each option by applying fundamental electromagnetic principles. R1V2 correctly identifies that while coil turns differ ($n_{1}>n_{2}$), the frequency of the induced alternating current depends on the rotational speed of the magnetic core rather than the number of turns: "Option B: The frequency of AC current produced by the two coils is equal; the frequency is determined by the rotational speed of the magnetic iron…" This analysis demonstrates R1V2’s ability to separate the relevant physical parameters (rotational frequency) from irrelevant ones (number of turns) when determining AC frequency, showcasing its understanding of electromagnetic induction principles. The model’s systematic elimination of incorrect options based on physical laws rather than pattern matching highlights its robust scientific reasoning capabilities.
Figure 6 showcases Skywork R1V2’s mathematical reasoning capabilities on a complex 3D geometry problem from the Chinese Gaokao. Faced with determining the volume of a sphere containing a square-based pyramid, R1V2 develops a structured solution approach: "My thought process might be: First step: Determine the relationship between the square base and the sphere… Second step: Consider the height of the pyramid…" The model strategically establishes a coordinate system, positioning the sphere’s center at the origin and the base in the xy-plane, then accurately applies the volume formula $(4/3)\pi R^{3}=(4/3)\cdot\pi\cdot 8=32\pi/3$. Particularly noteworthy is R1V2’s self-verification behavior, where it explicitly states "Let me check my steps again carefully to see if there are any errors…" before confirming its answer. This combination of systematic problem decomposition, precise mathematical calculation, and solution verification demonstrates the model’s sophisticated approach to mathematical reasoning tasks requiring spatial visualization.
@@ 4.4 Ablation Studies
As shown in Table 2, the SSB mechanism significantly improves model performance on key benchmarks. With our SSB approach, we achieved superior results on MathVista (74.0 vs. 73.8) while maintaining the strong performance on MMMU (73.6). More importantly, the proportion of effective samples remained above 60% with SSB, compared to below 40% without it. This efficiency gain demonstrates that SSB effectively addresses the vanishing advantages problem by ensuring a gradient-rich learning environment even as model responses converge.
@@
Full GRPO w/o SSB $<40\%$ 73.4 74.0
GRPO w/ SSB (10%) $>60\%$ 73.6 74.0
MPO vs. DPO vs. SFT
We evaluated our proposed Mixed Preference Optimization (MPO) approach against traditional Direct Preference Optimization (DPO) and standard Supervised Fine-Tuning (SFT). Table 3 presents the comparative results across multiple benchmarks and also reports hallucination rates. MPO consistently outperformed both alternatives across all metrics, with particularly notable improvements on reasoning-intensive tasks.
The hallucination rate using MPO was substantially lower (8.7%) compared to both DPO (12.6%) and SFT (18.4%). This reduction is especially significant as it indicates that our hybrid optimization approach effectively balances reasoning capability enhancement with factual accuracy. On complex mathematical reasoning tasks such as AIME 2024 and OlympiadBench, MPO achieved scores of 79.0 and 60.6 respectively, representing significant improvements over both SFT and DPO approaches. When combined with GRPO, our full approach achieves the best performance across all benchmarks (79.2 on AIME 2024 and 62.6 on OlympiadBench), confirming that our strategy successfully addresses the reasoning-generalization trade-off that has challenged previous models.
@@ 5 Conclusion and Future Work
The Selective Sample Buffer (SSB) mechanism effectively counters the "Vanishing Advantages" dilemma inherent in GRPO by maintaining a consistent supply of high-value training samples. Meanwhile, the Mixed Preference Optimization (MPO) strategy successfully balances reasoning specialization and generalization by integrating preference-based and rule-based reinforcement signals.

### v2

@@ 4.1.1 Text Reasoning Benchmarks
  • AIME 2024 [16]: This benchmark includes competition problems from the 2024 American Invitational Mathematics Examination (AIME), a prestigious and highly selective contest for elite high school students. It assesses advanced mathematical competencies, requiring deep conceptual understanding and rigorous logical reasoning skills.
@@
  • IFEVAL [45]: A benchmark designed to evaluate conditional reasoning in large language models, focusing on the models’ ability to correctly interpret and follow “if-then” statements. It tests for both factual accuracy and logical consistency in responding to hypothetical scenarios.
@@ 4.1.2 Multimodal Reasoning Benchmarks
Baselines We conduct comprehensive evaluations against several strong proprietary models, including Claude-3.5-Sonnet (20241022) [2], OpenAI-o4-mini [24], OpenAI-o1 [14], Gemini 2 Flash [32], and Kimi k1.5 longcot [34]. Additionally, we compare our method with advanced open-source models, such as Skywork-R1V1 [26], InternVL3-38B [46], QvQ-Preview-72B [35], Deepseek R1 [9] and Qwen2.5-VL-72B-Instruct [4]. For InternVL3-38B and other models where we report results not directly from their original papers, we conducted independent evaluations using our unified evaluation framework to ensure fair comparison.
@@ 4.2 Main Results
More impressively, R1V2 outperforms even larger models such as Qwen2.5-VL-72B (70.2% vs. 73.6% on MMMU) and QvQ-Preview-72B (70.3% vs. 73.6% on MMMU).
@@
Model MMMU Math- Math- Olympiad AIME LiveCode Live IFEVAL
Vista Vision Bench 24 bench Bench
Proprietary Models
Claude-3.5-Sonnet 70.4 67.7 - - - - - -
Gemini-2-Flash 70.7 73.1 41.3 - - - - -
Kimi-k1.5-longcot 70.0 74.9 53.3 - - - - -
OpenAI-o1 - - - - 74.3 63.4 72.2 -
OpenAI-o4-mini 81.6 84.3 58.0 - 93.4 74.6 78.1 -
@@
Skywork-R1V1 68.0 67.0 - - 72.0 57.2 54.6 72.5
DeepseekR1-671B - - - - 79.8 65.9 71.6 83.3
InternVL3-38B 70.1 75.1 34.2 - - - - -
Qwen2.5-VL-72B 70.2 74.8 38.1 40.4 - - - -
QvQ-Preview-72B 70.3 71.4 35.9 33.2 - - - -
Skywork-R1V2 (Ours) 73.6 74.0 49.0 62.6 78.9 63.6 73.2 82.9
@@
While R1V2 is optimized for reasoning capabilities, we also evaluated its performance on general vision understanding tasks to assess potential trade-offs. In document understanding, R1V2 maintains competitive performance (81.3% on AI2D [18], 79.0% on ChartQA [23] and TextVQA [30]) despite its reasoning focus, showing only a modest gap compared to specialized models. For video understanding, R1V2 achieves acceptable results (60.2% on Video-MME [11], 61.5% on MVBench [19]) but demonstrates stronger capabilities when temporal reasoning is required (1.92 on MMBench-Video [10] compared to InternVL2.5-38B’s 1.82).
Our analysis of hallucination tendencies revealed an important finding: while aggressive optimization for reasoning initially increased hallucination rates, our MPO approach effectively mitigated this issue, reducing hallucination from 18.4% with standard SFT to 8.7% with our full implementation. This balance allows R1V2 to maintain reasonable factual accuracy (68.9% on RealWorldQA [40]) while excelling at complex reasoning tasks. These results highlight the inherent trade-off between specialized reasoning and general visual understanding capabilities, with our current design deliberately prioritizing reasoning performance. Future versions will focus on enhancing general visual capabilities while preserving the strong reasoning foundation established in R1V2.
@@ 4.3 Qualitative Analysis
Figure 5 demonstrates R1V2’s physics reasoning capabilities through its analysis of an electromagnetic problem involving rotating magnetic fields and AC generators. When presented with a diagram showing two coils with different numbers of turns positioned perpendicular to each other, the model methodically evaluates each option by applying fundamental electromagnetic principles. R1V2 correctly identifies that while coil turns differ ($n_{1}>n_{2}$), the frequency of the induced alternating current depends on the rotational speed of the magnetic core rather than the number of turns: “Option B: The frequency of AC current produced by the two coils is equal; the frequency is determined by the rotational speed of the magnetic iron…” This analysis demonstrates R1V2’s ability to separate the relevant physical parameters (rotational frequency) from irrelevant ones (number of turns) when determining AC frequency, showcasing its understanding of electromagnetic induction principles. The model’s systematic elimination of incorrect options based on physical laws rather than pattern matching highlights its robust scientific reasoning capabilities.
Figure 6 showcases Skywork R1V2’s mathematical reasoning capabilities on a complex 3D geometry problem from the Chinese Gaokao. Faced with determining the volume of a sphere containing a square-based pyramid, R1V2 develops a structured solution approach: “My thought process might be: First step: Determine the relationship between the square base and the sphere… Second step: Consider the height of the pyramid…” The model strategically establishes a coordinate system, positioning the sphere’s center at the origin and the base in the xy-plane, then accurately applies the volume formula $(4/3)\pi R^{3}=(4/3)\cdot\pi\cdot 8=32\pi/3$. Particularly noteworthy is R1V2’s self-verification behavior, where it explicitly states “Let me check my steps again carefully to see if there are any errors…” before confirming its answer. This combination of systematic problem decomposition, precise mathematical calculation, and solution verification demonstrates the model’s sophisticated approach to mathematical reasoning tasks requiring spatial visualization.
@@ 4.4 Ablation Studies
As shown in Table 2, the SSB mechanism significantly improves model performance on key benchmarks. With our SSB approach, we achieved superior results on MMMU (73.6 vs. 73.4) while maintaining the strong performance on MathVista (74.0). More importantly, the proportion of effective samples remained above 60% with SSB, compared to below 40% without it. This efficiency gain demonstrates that SSB effectively addresses the “Vanishing Advantages” problem by ensuring a gradient-rich learning environment even as model responses converge.
@@
GRPO w/o SSB $<40\%$ 73.4 74.0
GRPO w/ SSB $>60\%$ 73.6 74.0
SFT vs MPO vs. Hybrid
We evaluated our proposed Hybrid approach [12] against Mixed Preference Optimization (MPO) and standard Supervised Fine-Tuning (SFT). Table 3 presents the comparative results across multiple benchmarks and also reports hallucination rates.
Although a specific configuration achieved peak MMMU performance (74.0), it incurred consistent slight performance degradation on other benchmarks, leading to its exclusion from the final release. The table shows that hallucination rate using MPO alone (8.7%) was substantially lower than both the hybrid MPO+GRPO approach (9.1%) and SFT (12.1%), demonstrating MPO’s strong standalone performance in reducing factual errors. On complex mathematical reasoning tasks, MPO achieved highest results on AIME 2024 (79.0) and competitive performance on OlympiadBench (60.6), significantly outperforming SFT (70.0 and 54.6, respectively). When combined with GRPO, our hybrid approach further improves generalization, achieving the highest scores on OlympiadBench (62.6) and near-optimal performance on AIME 2024(78.9). This confirms that hybrid approach successfully balances reasoning capability with generalization, addressing a key challenge faced by prior methods. Notably, SFT degraded model reasoning capabilities (as evidenced by its poor mathematical task performance), which motivated its complete removal from our subsequent training phases.
@@ 5 Conclusion and Future Work
Mixed Preference Optimization (MPO) strengthens the R1V model’s alignment via the Skywork-VL reward model, effectively reducing repetitive reasoning while maintaining robust generalization. To address the challenge of “Vanishing Advantages” in GRPO training, we introduce the Selective Sample Buffer (SSB), which strategically retains high-quality examples with distinct advantage signals, enabling stable policy updates via gradient-rich training samples.